OpenStudy (anonymous):

The height of a cylinder is 2 units. Its radius is 5 units. What is the volume of the cylinder, in cubic units? (Use π = 3.14) 157 cubic units 314 cubic units 426 cubic units 785 cubic units

Mehek (mehek14):

\(\tt{V=πr^2h}\)

OpenStudy (anonymous):

V= 3.14 5^2 2? Is that the right equation for that?

Mehek (mehek14):

yup \(v = 3.14 * 5^2*2\)

OpenStudy (anonymous):

3.14*5^2 = 78.5 8 2=157?

OpenStudy (anonymous):

Dont mind the 8 in the middle

Mehek (mehek14):

yup 157

OpenStudy (anonymous):

thank u !

Mehek (mehek14):

yw :)